# Towards Secure Virtual Elections: Multiparty Computation of Order Based Voting Rules

### v4

@@ 3.4 Verifying the legality of the cast ballots
Using the above observation regarding the linearity of secret sharing, then once the talliers receive $D^{\prime}$-out-of-$D$ shares in each entry in $\Gamma(Q)$, they can proceed to compute $D^{\prime}$-out-of-$D$ shares in the corresponding column sums, $Q_{m}$, $m\in[M]$, as we proceed to show.
In Copeland, conditions 2 and 3 in Theorem 3 imply that
$Q_{m}=\sum_{m^{\prime}\in[M]}Q(m^{\prime},m)=\sum_{m^{\prime}<m}Q(m^{\prime},m)-\sum_{m<m^{\prime}}Q(m,m^{\prime})\,.$ (5)
@@
Since the talliers hold shares in $Q(m^{\prime},m)$ for all $1\leq m^{\prime}<m\leq M$, they can use Eq. (5) and the linearity of secret sharing to compute shares in $Q_{m}$, $m\in[M]$. In Maximin, on the other hand, conditions 2 and 3 in Theorem 4 imply that
$Q_{m}=\sum_{m^{\prime}<m}Q(m^{\prime},m)-\sum_{m<m^{\prime}}Q(m,m^{\prime})+(M-m)\,.$ (6)
@@
Here too, the linearity of sharing and the relation in Eq. (6) enable the talliers to compute shares in $Q_{m}$, $m\in[M]$, also in the case of Maximin.
@@
$F(Q):=\left(\prod_{1\leq m^{\prime}<m\leq M}(Q_{m}-Q_{m^{\prime}})\right)^{2}\,.$ (7)
Condition 4, in both rules, holds iff $F(Q)\neq 0$. Hence, the talliers, who hold shares in $Q_{m}$, $m\in[M]$, may compute $F(Q)$ and then accept the ballot iff $F(Q)\neq 0$.
Privacy of the verification sub-protocol. A natural question that arises is whether the above described validation process poses a risk to the privacy of the voters. In other words, a voter that casts a legal ballot wants to be ascertained that the validation process only reveals that the ballot is legal, while all other information is kept hidden from the talliers. We proceed to examine that question.
The procedure for verifying condition 1 in Theorems 3 and 4 offers perfect privacy for honest voters. If the ballot $Q$ is legal then all multiplication gates will issue a zero output. Hence, apart from the legality of the ballot, the talliers will not learn anything about the content of the ballot.
The procedure for verifying condition 4 in Theorems 3 and 4 also offers perfect privacy, as we proceed to argue. If $Q$ is a valid ballot in Copeland, then the ordered tuple $(Q_{1},\ldots,Q_{M})$ is a permutation of the ordered tuple $(-M+1,-M+3,\ldots,M-3,M-1)$. This statement follows from the proof of condition 4 in Theorem 3. Hence, as can be readily verified,
$\prod_{1\leq m^{\prime}<m\leq M}(Q_{m}-Q_{m^{\prime}})=\pm 2^{\binom{M}{2}}\cdot\prod_{1\leq m^{\prime}<m\leq M}(m-m^{\prime})\,,$ (8)
where the sign of the product in Eq. (8) is determined by the signature of $(Q_{1},\ldots,Q_{M})$ when viewed as a permutation of the ordered tuple $(-M+1,-M+3,\ldots,M-3,M-1)$. Hence, as the talliers compute $F(Q)$, Eq. (7), which equals the square of the product on the left hand side of Eq. (8), then for any legal ballot they will always recover the same value, which is
$F(Q)=\left(2^{\binom{M}{2}}\cdot\prod_{1\leq m^{\prime}<m\leq M}(m-m^{\prime})\right)^{2}\,.$
@@
$F(Q)=\left(\prod_{1\leq m^{\prime}<m\leq M}(m-m^{\prime})\right)^{2}\,.$
Computational cost. Verifying condition 1 can be performed in parallel for all $M(M-1)/2$ entries in a given ballot, and also for several different ballots. Hence, in order to perform a batch validation of $B$ ballots, the talliers need to compute $BM(M-1)/2$ simultaneous multiplication gates.
The verification of condition 4 over a single ballot requires performing a sequence of $M(M-1)/2$ multiplications. Hence, in order to perform a batch validation of $B$ ballots, the talliers need to go through $M(M-1)/2$ rounds, where in each round they compute $B$ simultaneous multiplication gates.
@@ 3.5 Verifying the legality of the secret sharing
A malicious voter $V$ may attempt to sabotage the election by distributing to the $D$ talliers shares that do not correspond to a polynomial of degree $D^{\prime}-1$. Namely, if $\{s_{1},\ldots,s_{D}\}$ are the shares that $V$ distributed to $T_{1},\ldots,T_{D}$ in one of the entries in his ballot matrix, it is possible that there is no polynomial $g$ of degree (up to) $D^{\prime}-1$ such that $g(d)=s_{d}$ for all $d\in[D]$. By carefully selecting those shares, they may still pass the verification tests as described in Section 3.4 with some non-negligible probability, and then they would be integrated in the final computation of the winners. Since such shares do not correspond to a legal vote, they may sabotage the final computation of the winners (as we describe later on in Sections 3.6 and 3.7). To prevent such an attack, we explain herein how the talliers may detect it without learning anything on the submitted ballot (beyond the mere legality of the secret sharing that was applied to it). To that end the talliers proceed as follows:
  1. 1.
    Each tallier $T_{d}$, $d\in[D]$, produces a random number $r_{d}$ and distributes to all talliers $D^{\prime}$-out-of-$D$ shares in it, denoted $\{r_{d,1},\ldots,r_{d,D}\}$.
  2. 2.
    Each $T_{d}$, $d\in[D]$, computes $\hat{s}_{d}=s_{d}+\sum_{j\in[D]}r_{j,d}$. Namely, $T_{d}$ adds to the share that he had received from $V$ in the secret entry in $V$’s ballot matrix the shares received from all talliers in the random numbers that they had produced.
  3. 3.
    Each $T_{d}$, $d\in[D]$, broadcasts $\hat{s}_{d}$.
  4. 4.
    All talliers compute a polynomial $f$ of degree up to $D-1$ such that $f(d)=\hat{s}_{d}$, $d\in[D]$.
  5. 5.
    If $\deg f\leq D^{\prime}-1$ then the set of shares $\{s_{1},\ldots,s_{D}\}$ that $V$ had distributed is a legal $D^{\prime}$-out-of-$D$ sharing of some scalar in the field.
  6. 6.
    Otherwise, the ballot is rejected.
@@
The above procedure is correct and fully preserves the voter’s privacy.
We note that the first step can be executed even before the election period starts. Namely, once the number of registered voters, $N$, and the number of candidates, $M$, are determined, the talliers can produce $NM(M-1)/2$ random values and distribute shares in them to be used later on in masking the $M(M-1)/2$ entries in each voter’s ballot matrix.
3.6 Computing the winners in Copeland rule
The parameter $\alpha$ in Eq. (4) is always a rational number; typical settings of $\alpha$ are 00, $1$, or $\frac{1}{2}$  (Faliszewski et al., 2009). Assume that $\alpha=\frac{s}{t}$ for some integers $s$ and $t$. Then
$t\cdot\mbox{\bf w}(m)=t\cdot\sum_{m^{\prime}\neq m}1_{P(m,m^{\prime})>0}+s\cdot\sum_{m^{\prime}\neq m}1_{P(m,m^{\prime})=0}\,.$ (9)
@@ 3.6 Computing the winners in Copeland rule
$t\cdot\mbox{\bf w}(m)=$ $t\cdot\left\{\sum_{m^{\prime}>m}1_{P(m,m^{\prime})>0}+\sum_{m^{\prime}<m}1_{-P(m^{\prime},m)>0}\right\}+$ (10)
@@
In view of the above, the talliers may begin by computing secret shares in the bits of positivity in the first sum on the right-hand side of Eq. (10), by using the MPC sub-protocol described in Section 2.5. As for the bits of equality to zero in the second sum on the right-hand side of Eq. (10), the talliers can compute secret shares in them using the MPC sub-protocol described in Section 2.6. As the value of $t\cdot\mbox{\bf w}(m)$ is a linear combination of those bits, the talliers can then use the secret shares in those bits and Eq. (10) in order to get secret shares in $t\cdot\mbox{\bf w}(m)$, for each of the candidates, $C_{m}$, $m\in[M]$.
Next, they perform secure comparisons among the values $t\mbox{\bf w}(m)$, $m\in[M]$, in order to find the $K$ candidates with the highest scores. To do that, they need to perform $M-1$ secure comparisons (as described in Section 2.4) in order to find the candidate with the highest score, $M-2$ additional comparisons to find the next one, and so forth down to $M-K$ comparisons in order to find the $K$th winning candidate. Namely, the overall number of comparisons in this final stage is $\sum_{m=M-K}^{M-1}m=K\cdot\left(M-\frac{K+1}{2}\right)$, which is bounded by $M(M-1)/2$ for all $K<M$. Once this computational task is concluded, the talliers publish the indices of the $K$ winners (Line 11 in Protocol 1)).
We summarize the above described computation in Sub-protocol 2, which is an implementation of Line 11 in Protocol 1. It assumes that the talliers hold $D^{\prime}$-out-of-$D$ secret shares in $P(m,m^{\prime})$ for all $1\leq m<m^{\prime}\leq M$. Indeed, that computation has already taken place in Lines 9-10 of Protocol 1. Sub-protocol 2 starts with a computation of $D^{\prime}$-out-of-$D$ shares in all of the positivity bits and equality to zero bits that relate to the entries above the diagonal in $P$ (Lines 1-4). Then, in Lines 5-7, the talliers use those shares in order to obtain $D^{\prime}$-out-of-$D$ shares in $t\cdot\mbox{\bf w}(m)$ for each of the candidates, using Eq. (10); the $D^{\prime}$-out-of-$D$ shares in $t\cdot\mbox{\bf w}(m)$ are denoted $\{w_{d}(m):d\in[D]\}$. Finally, using the secure comparison sub-protocol, they find the $K$ winners (Lines 8-10).
Input: $T_{d}$, $d\in[D]$, has $G_{d}(m,m^{\prime})$ (a share in $P(m,m^{\prime})$) for all $1\leq m<m^{\prime}\leq M$.
@@
2       The talliers apply the positivity sub-protocol to translate $\{G_{d}(m,m^{\prime}):d\in[D]\}$ into shares $\{\sigma^{+}_{d}(m,m^{\prime}):d\in[D]\}$ in $1_{P(m,m^{\prime})>0}$; The talliers apply the positivity sub-protocol to translate $\{G_{d}(m,m^{\prime}):d\in[D]\}$ into shares $\{\sigma^{-}_{d}(m,m^{\prime}):d\in[D]\}$ in $1_{-P(m,m^{\prime})>0}$; The talliers apply the equality to zero sub-protocol to translate $\{G_{d}(m,m^{\prime}):d\in[D]\}$ into shares $\{\sigma^{0}_{d}(m,m^{\prime}):d\in[D]\}$ in $1_{P(m,m^{\prime})=0}$;
3forall $d\in[D]$ do
@@
6forall $k\in[K]:=\{1,\ldots,K\}$ do
7       The talliers perform $M-k$ invocations of the secure comparison sub-protocol over the $M-k+1$ candidates in C in order to find the $k$th elected candidate; The talliers output the candidate that was found and remove him from C;

### v7

@@ 3.4.2 Verifying condition 4
As the talliers received $D^{\prime}$-out-of-$D$ shares of each entry in $\Gamma(Q)$, they can compute $D^{\prime}$-out-of-$D$ shares of the corresponding row sums, $Q_{m}$, $m\in[M]$, as we proceed to show. In Copeland, conditions 2 and 3 in Theorem 1 imply that
$Q_{m}=\sum_{m^{\prime}\in[M]}Q(m,m^{\prime})=\sum_{m^{\prime}>m}Q(m,m^{\prime})-\sum_{m^{\prime}<m}Q(m^{\prime},m)\,.$ (4)
@@
Since the talliers hold shares of $Q(m^{\prime},m)$ for all $1\leq m^{\prime}<m\leq M$, they can use Eq. (4) and the linearity of secret sharing to compute shares of $Q_{m}$, $m\in[M]$. In Maximin, on the other hand, conditions 2 and 3 in Theorem 2 imply that
$Q_{m}=\sum_{m^{\prime}>m}Q(m,m^{\prime})-\sum_{m^{\prime}<m}Q(m^{\prime},m)+(m-1)\,.$ (5)
@@
Here too, the linearity of secret sharing and the relation in Eq. (5) enable the talliers to compute shares of $Q_{m}$, $m\in[M]$, also in the case of Maximin.
@@
$F(Q):=\prod_{1\leq m^{\prime}<m\leq M}(Q_{m}-Q_{m^{\prime}})\,.$ (6)
Condition 4, in both rules, holds iff $F(Q)\neq 0$. Hence, the talliers, who hold shares of $Q_{m}$, $m\in[M]$, may compute $F(Q)$ and then accept the ballot iff $F(Q)\neq 0$.
3.4.3 Privacy
A natural question that arises is whether the above described validation process poses a risk to the privacy of the voters. In other words, a voter that casts a legal ballot wants to be ascertained that the validation process only reveals that the ballot is legal, while all other information is kept hidden from the talliers. We proceed to examine that question.
The procedure for verifying condition 1 in Theorems 1 and 2 offers perfect privacy for honest voters. If the ballot $Q$ is legal then all computed values will be zero. Hence, apart from the legality of the ballot, the talliers will not learn anything about the content of the ballot.
The procedure for verifying condition 4 in Theorems 1 and 2 offers almost perfect privacy, in the following sense. If $Q$ is a valid ballot in Copeland, then the ordered tuple $(Q_{1},\ldots,Q_{M})$ is a permutation of the ordered tuple $(-M+1,-M+3,\ldots,M-3,M-1)$. This statement follows from the proof of condition 4 in Theorem 1. Hence, as can be readily verified, if $Q$ is a legal ballot then the value of $F(Q)$, Eq. (6), which the talliers compute in the validation procedure, equals
$F(Q)=\pm 2^{\binom{M}{2}}\cdot\prod_{1\leq m^{\prime}<m\leq M}(m-m^{\prime})\,,$ (7)
where the sign of the product in Eq. (7) is determined by the signature of $(Q_{1},\ldots,Q_{M})$ when viewed as a permutation of the ordered tuple $(-M+1,-M+3,\ldots,M-3,M-1)$. Hence, since a ballot in Copeland describes some ordering (permutation) of the candidates, the talliers will be able to infer the signature of that permutation, but nothing beyond that. To eliminate even that negligible leakage of information, the talliers will simply compute $F(Q)^{2}$ instead of $F(Q)$.
@@ 3.4.3 Privacy
$F(Q)=\pm\prod_{1\leq m^{\prime}<m\leq M}(m-m^{\prime})\,,$
where the sign of the above product is determined by the signature of $(Q_{1},\ldots,Q_{M})$ as a permutation of $(0,1,\ldots,M-2,M-1)$. Here too, the talliers will recover $F(Q)^{2}$ instead of $F(Q)$ to ensure perfect privacy.
@@ 3.5 Verifying the legality of the secret sharing
A malicious voter $V$ may attempt to sabotage the election by distributing to the $D$ talliers shares that do not correspond to a polynomial of degree $D^{\prime}-1$. Namely, if $\{s_{1},\ldots,s_{D}\}$ are the shares that $V$ distributed to $T_{1},\ldots,T_{D}$ in one of the entries in his ballot matrix, it is possible that there is no polynomial $g$ of degree (up to) $D^{\prime}-1$ such that $g(d)=s_{d}$ for all $d\in[D]$. By carefully selecting those shares, they may still pass the verification tests as described in Section 3.4 with some non-negligible probability, and then they would be integrated in the final computation of the winners. Since such shares do not correspond to a legal vote, they may sabotage the final computation of the winners (as we describe later on in Sections 3.6 and 3.7). To prevent such an attack, we explain herein how the talliers may detect it without learning anything on the submitted ballot (beyond the mere legality of the secret sharing that was applied to it).
Sub-protocol 2 performs the desired testing. The talliers apply it on the secret shares $\{s_{1},\ldots,s_{D}\}$ that they had received in any of the entries $s=Q(m^{\prime},m)$ in an incoming ballot matrix $Q$.
First, each tallier $T_{d}$, $d\in[D]$, produces a random number $r_{d}$ and distributes to all talliers $D^{\prime}$-out-of-$D$ shares of it (Lines 1-3). Then, each tallier $T_{d}$ adds to his share, $s_{d}$, the shares received from all talliers in the random numbers that they had produced, and broadcasts the result, denoted $\hat{s}_{d}$ (Lines 4-6). The talliers proceed to find, by means of interpolation, a polynomial $f$ that satisfies $f(d)=\hat{s}_{d}$ for all $d\in[D]$ (Line 7). If the degree of the polynomial is $D^{\prime}-1$ or less, the set of values $\{s_{1},\ldots,s_{D}\}$ constitute a legal $D^{\prime}$-out-of-$D$ sharing in some (still unknown) secret $s\in{\bf Z}_{p}$; otherwise, it is not (Lines 8-11).
Input: $T_{d}$, $d\in[D]$, has $s_{d}$, a supposedly $D^{\prime}$-out-of-$D$ share in some secret $s\in{\bf Z}_{p}$.
1
2forall $d\in[D]$ do
3
4      $T_{d}$ produces a random $r_{d}\in{\bf Z}_{p}$
5
6      $T_{d}$ distributes to all talliers $D^{\prime}$-out-of-$D$ shares of $r_{d}$, denoted $r_{d,1},\ldots,r_{d,D}$
7
8forall $d\in[D]$ do
9
10      $T_{d}$ computes $\hat{s}_{d}=s_{d}+\sum_{j\in[D]}r_{j,d}$
11
12      $T_{d}$ broadcasts $\hat{s}_{d}$
13
14All talliers compute a polynomial $f$ of degree up to $D-1$ such that $f(d)=\hat{s}_{d}$, $d\in[D]$
15
16if $\deg f\leq D^{\prime}-1$ then
17       Output ”A legal $D^{\prime}$-out-of-$D$ sharing”
18
19else
20       Output ”An illegal $D^{\prime}$-out-of-$D$ sharing”
21
Output : The legality of the secret sharing operation.
Sub-Protocol 2 Verifying that the shares distributed in some scalar secret constitute a legal $D^{\prime}$-out-of-$D$ sharing
@@ 20       Output ”An illegal $D^{\prime}$-out-of-$D$ sharing”
Sub-protocol 2 is correct and fully preserves the voter’s privacy.
Proof.
Assume that $V$ is honest. Then for each entry $s=Q(m^{\prime},m)$ in his ballot matrix there exists a polynomial $g$ of degree at most $D^{\prime}-1$ such that $s_{d}=g(d)$, $d\in[D]$. Define
$G:=g+\sum_{j\in[D]}g_{j}\,,$ (8)
where $g_{j}$, $j\in[D]$, is the polynomial of degree $D^{\prime}-1$ that $T_{j}$ used for generating the secret shares of its random value $r_{j}$ (Line 3 in Sub-protocol 2). Then
$G(d)=g(d)+\sum_{j\in[D]}g_{j}(d)=s_{d}+\sum_{j\in[D]}r_{j,d}=\hat{s}_{d}\,,\qquad d\in[D]\,.$ (9)
Hence, the interpolating polynomial $f$ that the talliers compute in Line 7 coincides with $G$. As $G$ is a sum of polynomials of degree $D^{\prime}-1$ at most, the validation in Lines 8-9 will pass successfully. In that case, the talliers would learn $G(0)=g(0)+\sum_{j\in[D]}r_{j}$. Here, $g(0)$ is the secret entry in $V$’s ballot matrix and $r_{j}$ is the secret random value that $T_{j}$ had chosen, $j\in[D]$. Hence, $G(0)$ reveals no information on $g(0)$. Since all other coefficients in $G$ are also random numbers that are not related to $g(0)$, this validation procedure provides perfect privacy for the voter.
Assume next that $V$ had distributed illegal shares for one of his ballot matrix entries. Namely, $V$ had distributed shares $\{s_{1},\ldots,s_{D}\}$ such that the minimum-degree polynomial $g$ that satisfies $g(d)=s_{d}$ for all $d\in[D]$ is of degree $t>D^{\prime}-1$. In that case, the polynomial $f$ that the talliers compute in Line 7 would be $f=g+\sum_{j\in[D]}g_{j}$. Since the degree of that polynomial is $t>D^{\prime}-1$, they would reject the ballot (Lines 10-11). Hence, the verification procedure is correct and provides perfect privacy. ∎
When the talliers receive shares in some ballot matrix, they must first execute Sub-protocol 2 for each of its entries, and only if that validation passes successfully they will proceed to perform the validation described in Section 3.4.
Finally, we note that Lines 1-3 in Sub-protocol 2 can be executed even before the election period starts. Namely, once the number of registered voters, $N$, and the number of candidates, $M$, are determined, the talliers can produce $NM(M-1)/2$ random values and distribute shares of them to be used later on in masking the $M(M-1)/2$ entries in each voter’s ballot matrix.
3.6 An MPC computation of the winners in the Copeland rule
The parameter $\alpha$ in Eq. (2) is always a rational number; typical settings of $\alpha$ are 00, $1$, or $\frac{1}{2}$  [25]. Assume that $\alpha=\frac{s}{t}$ for some integers $s$ and $t$. Then
$t\cdot\mbox{\bf w}(m)=t\cdot\sum_{m^{\prime}\in[M]\setminus\{m\}}1_{P(m,m^{\prime})>0}+s\cdot\sum_{m^{\prime}\in[M]\setminus\{m\}}1_{P(m,m^{\prime})=0}\,.$ (10)
@@ 3.6 An MPC computation of the winners in the Copeland rule
$t\cdot\mbox{\bf w}(m)$ $=$ (11)
$t\cdot\left\{\sum_{m^{\prime}>m}1_{P(m,m^{\prime})>0}+\sum_{m^{\prime}<m}1_{-P(m^{\prime},m)>0}\right\}+$
@@
In view of the above, the talliers may begin by computing secret shares of the bits of positivity in the first sum on the right-hand side of Eq. (11), by using the MPC sub-protocol described in Section A.3. As for the bits of equality to zero in the second sum on the right-hand side of Eq. (11), the talliers can compute secret shares of them using the MPC sub-protocol described in Section A.4. As the value of $t\cdot\mbox{\bf w}(m)$ is a linear combination of those bits, the talliers can then use the secret shares of those bits and Eq. (11) in order to get secret shares of $t\cdot\mbox{\bf w}(m)$, for each of the candidates, $C_{m}$, $m\in[M]$.
Next, they perform secure comparisons among the values $t\mbox{\bf w}(m)$, $m\in[M]$, in order to find the $K$ candidates with the highest scores. To do that, they need to perform $M-1$ secure comparisons (as described in Section 2.2 and then further discussed in Section A.2) in order to find the candidate with the highest score, $M-2$ additional comparisons to find the next one, and so forth down to $M-K$ comparisons in order to find the $K$th winning candidate. Namely, the overall number of comparisons in this final stage is
$\sum_{m=M-K}^{M-1}m=K\cdot\left(M-\frac{K+1}{2}\right)\,.$
The above number is bounded by $M(M-1)/2$ for all $K<M$. Once this computational task is concluded, the talliers publish the indices of the $K$ winners (Line 11 in Protocol 1)).
We summarize the above described computation in Sub-protocol 3, which is an implementation of Line 11 in Protocol 1. It assumes that the talliers hold $D^{\prime}$-out-of-$D$ secret shares of $P(m,m^{\prime})$ for all $1\leq m<m^{\prime}\leq M$. Indeed, that computation has already taken place in Lines 9-10 of Protocol 1. Sub-protocol 3 starts with a computation of $D^{\prime}$-out-of-$D$ shares of all of the positivity bits and equality to zero bits that relate to the entries above the diagonal in $P$ (Lines 1-4). Then, in Lines 5-7, the talliers use those shares in order to obtain $D^{\prime}$-out-of-$D$ shares of $t\cdot\mbox{\bf w}(m)$ for each of the candidates, using Eq. (11); the $D^{\prime}$-out-of-$D$ shares of $t\cdot\mbox{\bf w}(m)$ are denoted $\{w_{d}(m):d\in[D]\}$. Finally, using the secure comparison sub-protocol, they find the $K$ winners (Lines 8-10).
Input: $T_{d}$, $d\in[D]$, has $G_{d}(m,m^{\prime})$ (a share of $P(m,m^{\prime})$) for all $1\leq m<m^{\prime}\leq M$.
@@
2       The talliers apply the positivity sub-protocol to translate $\{G_{d}(m,m^{\prime}):d\in[D]\}$ into shares $\{\sigma^{+}_{d}(m,m^{\prime}):d\in[D]\}$ in $1_{P(m,m^{\prime})>0}$;
3      The talliers apply the positivity sub-protocol to translate $\{G_{d}(m,m^{\prime}):d\in[D]\}$ into shares $\{\sigma^{-}_{d}(m,m^{\prime}):d\in[D]\}$ in $1_{-P(m,m^{\prime})>0}$;
4      The talliers apply the equality to zero sub-protocol to translate $\{G_{d}(m,m^{\prime}):d\in[D]\}$ into shares $\{\sigma^{0}_{d}(m,m^{\prime}):d\in[D]\}$ in $1_{P(m,m^{\prime})=0}$;
5forall $d\in[D]$ do
@@
8forall $k\in[K]:=\{1,\ldots,K\}$ do
9       The talliers perform $M-k$ invocations of the secure comparison sub-protocol over the $M-k+1$ candidates in C in order to find the $k$th elected candidate;
10      The talliers output the candidate that was found and remove him from C;